OpenStudy (anonymous):

In a population of plants, the allele for long stems (S) is completely dominant over the allele for short stems (s). If 35% of the population has short stems, calculate the percentage of the population that is expected to be heterozygous (Ss) 48 percent 59 percent 41 percent 17 percent

OpenStudy (anonymous):

PLEASE HELP <3 :)

OpenStudy (aaronq):

we need to use p+q=1 and \(p^2+2pq+q^2\) we say that S=p and s=q short stems are \(ss\), so \(q^2\)=35%=0.35 then q=\(\sqrt{0.35}=0.592\) using the original equation p+q=1, p=1-0.592=0.408 so, the heterozygotes are 2pq=2(0.408)(0.592)=0.483=48.3% are heterozygotes

OpenStudy (anonymous):

Its 48